# Depression Detection and Analysis using Large Language Models on Textual and Audio-Visual Modalities

### v1

@@ II-C Large Language Models Based Approaches
Furthermore, to demonstrate the capabilities of large language models, A. Anand et al.[30] finetuned an LLM through instructional calibration on a dataset tailored to High School Physics and utilizing retrieval augmentation. Their finetuned retrieval-augmented model, SciPhy-RAG, demonstrates significant improvements over Vicuna-7b.
Also, in their other papers [31] and [32] A. Anand et al. propose MM-PhyQA, a dataset featuring high school-level multimodal physics problems and ’Mathquest’ a maths dataset derived from 11th and 12th standard NCERT Mathematics textbooks respectively. These papers show their study to evaluate the performance of LLMs on these datasets, and their approach highlights the potential for enhancing LLMs’ capability in textual corpora with specialized datasets and techniques.
Moreover, using the same dataset MM-PhyQA, [33] proposed an LLM-based chatbot to answer multimodal physics multiple-choice questions (MCQs). They demonstrated that their Reinforcement Learning from Human Feedback (RLHF) and Image Captioning techniques improve the quality and accuracy of the model’s responses compared to supervised fine-tuned LLMs.

### v2

@@ II-C Large Language Models Based Approaches
Furthermore, to demonstrate the capabilities of large language models, their applications can be seen across various domains, particularly in the scientific and educational sectors, where they have demonstrated significant potential. In the scientific domain, LLMs are proving beneficial for tasks such as citation generation [44, 49, 55], where they assist in creating precise and contextually relevant citations, and grammatical error correction [46, 47], helping improve the fluency and accuracy of scholarly writing. These advancements have made LLMs valuable tools in enhancing the quality of scientific communication.
In the educational domain, the impact of LLMs extends to disciplines such as mathematics [50, 51, 52, 53] and physics [45, 48], where LLMs are being used to develop innovative methodologies that enhance both teaching and learning experiences. For example, in the context of physics, retrieval augmentation and fine-tuning of LLMs have been applied to improve the accuracy of physics-related question-answering systems. Additionally, LLMs are playing a role in improving learner engagement by assessing attention in e-classrooms, with frameworks like ExCEDA [54] and ECLIPSE Dataset [56] being developed to unlock new attention paradigms. Furthermore, LLMs are contributing to the field of code generation [57], where they help automate the creation of programming code, making it easier for students and researchers to engage in computational tasks. These applications reflect the versatility of LLMs, effectively bridging their capabilities with the specific needs of both scientific research and educational development.